# Geometric structures on the quaternionic unit ball and slice regular Möbius transformations

### v1

@@ 4. Hermitian structure on $\mathbb{D}$
(4.3) $h_{z}(\alpha,\beta)=\operatorname{d\!}\varphi_{z}(\alpha)\overline{\operatorname{d\!}\varphi_{z}(\beta)},$
for every $\alpha,\beta\in\mathbb{C}$. If $\psi$ is some other Möbius transformation of $\mathbb{D}$ mapping $\psi(z)=0$, then the transformation $\varphi\circ\psi^{-1}$ fixes 00 and so there exists $(a,d)\in\mathbb{T}\times\mathbb{T}$ such that $\varphi\circ\psi^{-1}=R_{(a,d)}$, where $R_{(a,d)}$ stands for the right action of the element $(a,d)\in\mathbb{T}\times\mathbb{T}$ over $\mathbb{D}$. We can rewrite this as
@@
$\operatorname{d\!}\varphi_{z}(\alpha)\overline{\operatorname{d\!}\varphi_{z}(\beta)}=\operatorname{d\!}R_{(a,d)}(\operatorname{d\!}\psi_{z}(\alpha))\overline{\operatorname{d\!}R_{(a,d)}(\operatorname{d\!}\psi_{z}(\beta))}=\operatorname{d\!}\psi_{z}(\alpha)\overline{\operatorname{d\!}\psi_{z}(\beta)},$
@@
The expression follows by an explicit computation of $\operatorname{d\!}\varphi_{z}$ for any choice of $\varphi$ as in the previous discussion. The $\mathrm{U}(1,1)$-invariance can be proved easily from its definition and the fact that $\mathrm{U}(1,1)$ acts on $\mathbb{D}$. ∎
@@ 5. Slice geometric structures on $\mathbb{B}$
which are defined for $\alpha,\beta\in\mathbb{H}$ as well. In this case, $G_{0}$ is the canonical positive definite inner product of $\mathbb{H}\simeq\mathbb{R}^{4}$ and $\Omega_{0}$ is an $\operatorname{Im}(\mathbb{H})$-valued anti-symmetric $\mathbb{R}$-bilinear form. We will think of these three forms, $H_{0}$, $G_{0}$ and $\Omega_{0}$, as defined on the tangent space to $\mathbb{B}$ at the origin 00. Note that equation (5.1) corresponds to the classical complex case given by equation (4.1). Similarly, equations (5.2) correspond to the real and imaginary parts of the latter.
@@
    There does not exist $\mathbb{H}$-valued left Hermitian positive definite tensor forms on $\mathbb{B}$ that are also $\mathrm{Sp}(1,1)$-invariant.
@@
    There does not exist $\operatorname{Im}(\mathbb{H})$-valued $2$-forms on $\mathbb{B}$ whose value at 00 is $\Omega_{0}$ that are also $\mathrm{Sp}(1,1)$-invariant.
@@
It is well known (see [2, 12]) that the $\mathrm{Sp}(1,1)$-action on $\mathbb{B}$ is transitive thus realizing the latter as a homogeneous space
$\mathbb{B}=(\mathrm{Sp}(1)\times\mathrm{Sp}(1))\backslash\mathrm{Sp}(1,1).$
Furthermore, $(\mathrm{Sp}(1,1),\mathrm{Sp}(1)\times\mathrm{Sp}(1))$ is a Riemannian symmetric pair with $\mathrm{Sp}(1,1)$ a non-compact simple Lie group. Then, the first claim of (1) follows from the theory of Riemannian symmetric spaces. We refer to [12] for definitions and details. On the other hand, for every $u,v\in\mathrm{Sp}(1)$ and $\alpha,\beta\in\mathbb{H}$ we have
@@
thus proving that $G_{0}$ is $\mathrm{Sp}(1)\times\mathrm{Sp}(1)$-invariant. Hence, the second claim in (1) follows from the theory of Riemannian symmetric spaces as well. Finally, the expression for $\widehat{G}$ follows from the results found in [2]. Next, we claim that $\Omega_{0}$ is not $\mathrm{Sp}(1)\times\mathrm{Sp}(1)$-invariant, and so that the same holds for $H_{0}$. To see this, let us assume that $u\in\mathrm{Sp}(1)$ satisfies $\operatorname{Im}(\overline{u}\alpha\overline{\overline{u}\beta})=\operatorname{Im}(\alpha\overline{\beta})$, for all $\alpha,\beta\in\mathbb{H}$. Then, by choosing $\beta=1$ we conclude that $\operatorname{Im}(\overline{u}\alpha u)=\operatorname{Im}(\alpha)$, which is easily seen to imply that $\alpha u=u\alpha$, for all $\alpha\in\mathbb{H}$, because $|u|=1$. Hence, we conclude that $u=\pm 1$, thus proving our claim. Let us assume the existence of an $\mathrm{Sp}(1,1)$-invariant $\mathbb{H}$-valued left Hermitian positive definite tensor form on $\mathbb{B}$ and let $T_{0}$ be its value at 00. In particular, $T_{0}$ is an $\mathrm{Sp}(1)\times\mathrm{Sp}(1)$-invariant $\mathbb{H}$-valued left Hermitian positive definite form on $\mathbb{H}$. By the classification of the latter (see [11, Chapter 2]), there exists $c>0$ such that $H_{0}=cT_{0}$. This implies that $H_{0}$ is $\mathrm{Sp}(1)\times\mathrm{Sp}(1)$-invariant which is a contradiction. This proves (2). Finally, the non-invariance of $\Omega_{0}$ for the $\mathrm{Sp}(1)\times\mathrm{Sp}(1)$-action and arguments similar to those used above yield (3). ∎
@@
Theorem 5.1 can be seen as proving that non-regular Möbius transformations do not provide invariant Hermitian and symplectic geometric structures on $\mathbb{B}$ of a quaternionic nature. By comparing the proof of Theorem 5.1 with the constructions from section 4 one can readily see that this fact is a direct consequence of the non-commutativity of $\mathbb{H}$. On the other hand, Theorem 5.1(1) is a restatement of the well known existence and uniqueness of an $\mathrm{Sp}(1,1)$-invariant metric on $\mathbb{B}$ (see [2, 12]). In fact, the Riemannian metric $\widehat{G}$ above is indeed the one studied in [2].
In order to obtain Hermitian and symplectic-like geometric structures on $\mathbb{B}$ of a true quaternionic nature we set our sights on $\mathcal{M}(\mathbb{B})$, the regular Möbius transformations. Following the ideas from section 4 we first note that, in contrast to the arguments found in the proof of Theorem 5.1, the tensor $H_{0}$, and so $G_{0}$ and $\Omega_{0}$, is invariant under the regular Möbius transformations that fix the origin.
@@
Let us denote by $\mathcal{M}_{0}(\mathbb{B})$ the set of regular Möbius transformations of $\mathbb{B}$ that fix 00. Then, the following properties hold.
@@
    yields maps of $\mathbb{B}$ that realize $\mathcal{M}_{0}(\mathbb{B})$. In other words, $\mathcal{M}_{0}(\mathbb{B})$ consists of the set of transformations $R_{u}$, for $u\in\mathrm{Sp}(1)$. In particular, $\mathcal{M}_{0}(\mathbb{B})$ is a group isomorphic to $\mathrm{Sp}(1)$ acting on the right on $\mathbb{B}$ and fixing 00.
@@
    The differential at 00 of the $\mathcal{M}_{0}(\mathbb{B})$-action on $\mathbb{B}$ leaves invariant the form $H_{0}$. In particular, the same holds for $G_{0}$ and $\Omega_{0}$.
@@
Clearly, for every $u\in\mathrm{Sp}(1)$, the map $R_{u}$ is a regular Möbius transformation fixing 00. Conversely, let us assume that $\mathcal{F}$ is a regular Möbius transformation of $\mathbb{B}$ that fixes 00. By Proposition 3.2, there exist $a\in\mathbb{B}$ and $u\in\mathrm{Sp}(1)$ such that $\mathcal{F}=R_{u}\circ\mathcal{F}_{a}$, where we have used the notation from Proposition 3.3. By assumption we have
$0=\mathcal{F}(0)=\mathcal{F}_{a}(0)u=au$
where we have used Proposition 3.3(1). This implies $a=0$ and so that $\mathcal{F}(q)=\mathcal{F}_{0}(q)u=q(-u)$, for all $q\in\mathbb{B}$. Hence, we have obtained the required description of $\mathcal{M}_{0}(\mathbb{B})$. The rest of the claims from (1) are now immediate. Since the $\mathcal{M}_{0}(\mathbb{B})$-action on $\mathbb{B}$ is $\mathbb{R}$-linear, its differential has the same expression. Hence, the trivial identity
@@
The $\mathrm{Sp}(1)$-invariance of $H_{0}$ is the first step to build geometric structures on $\mathbb{B}$ out of regular Möbius transformations. It is important to note that Proposition 5.3 has provided a sort of isotropy at 00 for such transformations, which turns out to be an actual group. However, for the composition of maps, $\mathcal{M}(\mathbb{B})$ is not a group and so at this point the full group theoretic argument used in section 4 for the classical complex case breaks down. Nevertheless, the next result provides the analog of equation (4.4) for regular Möbius transformations.
@@
$H_{q}(\alpha,\beta)$ $=\operatorname{d\!}\mathcal{F}_{q}(\alpha)\overline{\operatorname{d\!}\mathcal{F}_{q}(\beta)}$
$G_{q}(\alpha,\beta)$ $=\operatorname{Re}\big{(}\operatorname{d\!}\mathcal{F}_{q}(\alpha)\overline{\operatorname{d\!}\mathcal{F}_{q}(\beta)}\big{)}$
$\Omega_{q}(\alpha,\beta)$ $=\operatorname{Im}\big{(}\operatorname{d\!}\mathcal{F}_{q}(\alpha)\overline{\operatorname{d\!}\mathcal{F}_{q}(\beta)}\big{)}$
@@
$H_{q}(\alpha,\beta)=\frac{(1-q^{2})^{-1}(\alpha-q\alpha q)(\overline{q-q\beta q})(1-\overline{q}^{2})^{-1}}{(1-|q|^{2})^{2}},$
@@
$\operatorname{d\!}\>(\mathcal{F}_{1})_{q}(\alpha)\overline{\operatorname{d\!}\>(\mathcal{F}_{1})_{q}(\beta)}=\big{(}\operatorname{d\!}\>(\mathcal{F}_{2})_{q}(\alpha)u\big{)}\overline{\big{(}\operatorname{d\!}\>(\mathcal{F}_{2})_{q}(\beta)u\big{)}}=\operatorname{d\!}\>(\mathcal{F}_{2})_{q}(\alpha)\overline{\operatorname{d\!}\>(\mathcal{F}_{2})_{q}(\beta)}.$
This proves that all three tensors $H$, $G$ and $\Omega$ are well defined. Hence, the definition of $H$ clearly implies claims (1) and (2). For a given $q\in\mathbb{B}$, it is straightforward to compute the formula for $H_{q}$ by choosing $\mathcal{F}_{q}$ as the transformation mapping $q$ to 00 and using the expression for its differential at $q$ obtained in Proposition 3.3(2). The formula for $H$ now implies the smoothness of all three tensors. Finally, to prove the positive definiteness of $H$ let us choose $q\in\mathbb{B}$ and $\alpha\in\mathbb{H}$. Then, we clearly have
@@ 6. Properties of the slice geometric structures on $\mathbb{B}$
The tensors $G$ and $\Omega$ can be given explicit expressions by simply taking the real and imaginary parts, respectively, of $H$. However, we can provide some further simplified formulas. To achieve this, we introduce some notation.
@@
where $J$ is any other element of $\mathbb{S}$ that anti-commutes with $I$. This decomposition is orthogonal for the canonical inner product of $\mathbb{H}\simeq\mathbb{R}^{4}$. In particular, $\mathbb{C}_{I}J$ is the orthogonal complement for such inner product and so it does not depend on the choice of $J$. We will denote from now on by $\pi_{I}$ and $\pi_{I}^{\perp}$ the orthogonal projections from $\mathbb{H}$ onto $\mathbb{C}_{I}$ and $(\mathbb{C}_{I})^{\perp}=\mathbb{C}_{I}J$, respectively.
It follows from the proof of Theorem 5.7 that $G$ is indeed a smooth Riemannian metric. The next result further provides for $G$ a pair of useful simplified formulas.
@@
$G_{q}(\alpha,\beta)$ $=\frac{\operatorname{Re}(\alpha\overline{\beta})}{(1-|q|^{2})^{2}}-4\frac{\operatorname{Im}(q)^{2}\operatorname{Re}(\pi_{I}^{\perp}(\alpha)\pi_{I}^{\perp}(\beta))}{|1-q^{2}|^{2}(1-|q|^{2})^{2}}$
$=\widehat{G}_{q}(\alpha,\beta)-4\frac{\operatorname{Im}(q)^{2}\operatorname{Re}(\pi_{I}^{\perp}(\alpha)\pi_{I}^{\perp}(\beta))}{|1-q^{2}|^{2}(1-|q|^{2})^{2}},$
for every $q\in\mathbb{D}_{I}$, where $I\in\mathbb{S}$, and $\alpha,\beta\in\mathbb{H}$. Recall that $\widehat{G}$ is the $\mathrm{Sp}(1,1)$-invariant metric on $\mathbb{B}$ from Theorem 5.1(1).
@@
Since $\operatorname{Re}(\alpha\beta)=\operatorname{Re}(\beta\alpha)$ for every $\alpha,\beta\in\mathbb{H}$, the first expression follows from Theorem 5.7. We also note that for $q\in\mathbb{R}$, the second and third expressions for $G_{q}$ are obvious. Hence, we proceed to manipulate the expression
@@
under the assumption that $q\in\mathbb{C}_{I}\setminus\mathbb{R}$ for some $I\in\mathbb{S}$. We first observe that
$\operatorname{Re}((\alpha-q\alpha q)(\overline{\beta-q\beta q}))=(1+|q|^{4})\operatorname{Re}(\alpha\overline{\beta})-\operatorname{Re}(q\alpha q\overline{\beta})-\operatorname{Re}(\overline{q}\alpha\overline{q}\overline{\beta}).$
Let us choose $J\in\mathbb{S}$ that anti-commutes with $I$. If we write $\alpha=\alpha_{1}+\alpha_{2}J$ and $\beta=\beta_{1}+\beta_{2}J$, with $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}\in\mathbb{C}_{I}$, then a direct computation yields
@@
$=(1-2\operatorname{Re}(q^{2})+|q|^{4})\operatorname{Re}(\alpha\overline{\beta})+2(\operatorname{Re}(q^{2})-|q|^{2})\operatorname{Re}(\alpha_{2}\overline{\beta_{2}})$
$=|1-q^{2}|^{2}\operatorname{Re}(\alpha\overline{\beta})-4\operatorname{Im}(q)^{2}\operatorname{Re}(\alpha_{2}\overline{\beta_{2}}),$
@@
The third formula now follows from Theorem 5.1(1). ∎
As a consequence, we conclude that the Riemannian geometry on $\mathbb{D}_{I}=\mathbb{B}\cap\mathbb{C}_{I}$ induced from the slice Riemannian metric $G$ of $\mathbb{B}$ is exactly the usual hyperbolic geometry. More precisely, we have the next result. We will denote by $g_{I}$ the Riemannian metric on $\mathbb{D}_{I}$ induced by the canonical identification $\mathbb{D}_{I}\simeq\mathbb{D}$.

### v2

@@ 4. Hermitian structure on $\mathbb{D}$
[…]
for every $\alpha,\beta\in\mathbb{C}$. If $\psi$ is some other Möbius transformation of $\mathbb{D}$ mapping $\psi(z)=0$, then the transformation $\varphi\circ\psi^{-1}$ fixes 00 and so we conclude the existence of $(a,d)\in\mathbb{T}\times\mathbb{T}$ such that $\varphi\circ\psi^{-1}=R_{(a,d)}$, where $R_{(a,d)}$ stands for the right action of the element $(a,d)\in\mathbb{T}\times\mathbb{T}$ over $\mathbb{D}$. Hence, we can rewrite this as
@@
[…]
@@
[…]
@@ 5. Slice geometric structures on $\mathbb{B}$
which are defined for $\alpha,\beta\in\mathbb{H}$ as well. In this case, $G_{0}$ is the canonical positive definite inner product of $\mathbb{H}\simeq\mathbb{R}^{4}$ and $\Omega_{0}$ is an $\operatorname{Im}(\mathbb{H})$-valued anti-symmetric $\mathbb{R}$-bilinear form. We will think of these three forms, $H_{0}$, $G_{0}$ and $\Omega_{0}$, as defined on the tangent space to $\mathbb{B}$ at the origin 00. Note that equation (5.1) corresponds to the classical complex case given by equation (4.1). Similarly, the forms (5.2) correspond to the real part and $i$-times the imaginary part of (4.1). To clarify the point, it is important to note that $\Omega_{0}$ is vector-valued, while the imaginary part of (4.1) is real-valued.
@@
    There do not exist $\mathbb{H}$-valued left Hermitian positive definite tensor forms on $\mathbb{B}$ that are also $\mathrm{Sp}(1,1)$-invariant.
@@
    There do not exist $\operatorname{Im}(\mathbb{H})$-valued $2$-forms on $\mathbb{B}$ whose value at 00 is $\Omega_{0}$ that are also $\mathrm{Sp}(1,1)$-invariant.
@@
It is well known (see [3, 15]) that the right $\mathrm{Sp}(1,1)$-action on $\mathbb{B}$ is transitive. Furthermore, if we consider (as before) $\mathrm{Sp}(1)\times\mathrm{Sp}(1)$ as a diagonally embedded subgroup of $\mathrm{Sp}(1,1)$, then such subgroup is precisely the stabilizer of 00 for the $\mathrm{Sp}(1,1)$-action. This can be used to obtain a realization of $\mathbb{B}$ as a homogeneous space that can be described as follows. Our main reference is [15] which considers left actions as opposed to our right actions. However, the standard correspondence between these two types of actions can be used to apply the results in [15] to our setup.
We recall that the coset space
$(\mathrm{Sp}(1)\times\mathrm{Sp}(1))\backslash\mathrm{Sp}(1,1)$
admits a unique manifold structure such that the $\mathrm{Sp}(1,1)$-action on the right is smooth. This is a consequence of [15, Theorem 4.2, Chapter II]. If we consider the map given by
$(\mathrm{Sp}(1)\times\mathrm{Sp}(1))\backslash\mathrm{Sp}(1,1)$ $\rightarrow\mathbb{B}$
$(\mathrm{Sp}(1)\times\mathrm{Sp}(1))A$ $\mapsto F_{A}(0)=0\cdot A,$
then the previous remarks on the $\mathrm{Sp}(1,1)$-action show that this is a well-defined bijection. Furthermore, Theorem 3.2 and Proposition 4.3 in Chapter II of [15] imply that such map is in fact a diffeomorphism, thus providing a homogeneous space realization of $\mathbb{B}$. We also note that such diffeomorphism is $\mathrm{Sp}(1,1)$-equivariant for the right actions on both domain and target.
On the other hand, by [15, Table V, page 518], the pair
$(\mathrm{Sp}(1,1),\mathrm{Sp}(1)\times\mathrm{Sp}(1))$
is a so-called Riemannian symmetric pair (see Definition in [15, page 209]) of type CII. Hence, [15, Proposition 3.4, Chapter IV] implies the existence of a $\mathrm{Sp}(1,1)$-invariant Riemannian metric on the quotient space
$(\mathrm{Sp}(1)\times\mathrm{Sp}(1))\backslash\mathrm{Sp}(1,1),$
which thus translates into a $\mathrm{Sp}(1,1)$-invariant Riemannian metric on $\mathbb{B}$ as a consequence of the equivariance satisfied by the diffeomorphism between these two manifolds. This proves the existence of the Riemannian metric as stated in (1). To establish uniqueness, we recall that $\mathrm{Sp}(1,1)$ is a simple Lie group (see Definition in [15, page 131]) and that the $\mathrm{Sp}(1)\times\mathrm{Sp}(1)$-action on the tangent space at 00 in $\mathbb{B}$ is irreducible; this follows from the fact that the symmetric pair $(\mathrm{Sp}(1,1),\mathrm{Sp}(1)\times\mathrm{Sp}(1))$ is listed in [15, Table V, page 518]. Hence, by the remarks at the end of page 255 in [15] the Riemannian metric in $\mathbb{B}$ is unique, up to a constant positive factor. Also note that, for every $u,v\in\mathrm{Sp}(1)$ and $\alpha,\beta\in\mathbb{H}$ we have
@@
thus proving that $G_{0}$ is $\mathrm{Sp}(1)\times\mathrm{Sp}(1)$-invariant, and so $G_{0}$ is unique, up to a constant factor, by the irreducibility mentioned above. To complete the proof of (1) we note that the expression for $\widehat{G}$ has been computed in [3]: for further details see [3, Theorem 1.9] and the remarks before this result.
Next, we claim that $\Omega_{0}$ is not $\mathrm{Sp}(1)\times\mathrm{Sp}(1)$-invariant, and so that the same holds for $H_{0}$. To see this, let us assume that $u\in\mathrm{Sp}(1)$ satisfies $\operatorname{Im}(\overline{u}\alpha\overline{\overline{u}\beta})=\operatorname{Im}(\alpha\overline{\beta})$, for all $\alpha,\beta\in\mathbb{H}$. Then, by choosing $\beta=1$ we conclude that $\operatorname{Im}(\overline{u}\alpha u)=\operatorname{Im}(\alpha)$, which is easily seen to imply that $\alpha u=u\alpha$, for all $\alpha\in\mathbb{H}$, because $|u|=1$. Hence, we conclude that $u=\pm 1$, thus proving our claim.
Let us assume the existence of an $\mathrm{Sp}(1,1)$-invariant $\mathbb{H}$-valued left Hermitian positive definite tensor form on $\mathbb{B}$ and let $T_{0}$ be its value at 00. In particular, $T_{0}$ is an $\mathrm{Sp}(1)\times\mathrm{Sp}(1)$-invariant $\mathbb{H}$-valued left Hermitian positive definite form on $\mathbb{H}$. By The Basis Theorem 2.46 from [14], applied to the $\mathbb{H}$-Hermitian symmetric type (see [14, page 21]), there exists $c>0$ such that $H_{0}=cT_{0}$. This implies that $H_{0}$ is $\mathrm{Sp}(1)\times\mathrm{Sp}(1)$-invariant, which is a contradiction. This proves (2).
Finally, the non-invariance of $\Omega_{0}$ for the $\mathrm{Sp}(1)\times\mathrm{Sp}(1)$-action and arguments similar to those used above yield (3). ∎
@@
Theorem 5.1 can be seen as proving that non-regular Möbius transformations do not provide invariant Hermitian and symplectic geometric structures on $\mathbb{B}$ of a quaternionic nature. By comparing the proof of Theorem 5.1 with the constructions from section 4 one can readily see that this fact is a direct consequence of the non-commutativity of $\mathbb{H}$.
On the other hand, Theorem 5.1(1) is a restatement of the well known existence and uniqueness of an $\mathrm{Sp}(1,1)$-invariant metric on $\mathbb{B}$ (see [3, 15]). In fact, the Riemannian metric $\widehat{G}$ above is indeed the one studied in [3].
In order to obtain Hermitian and symplectic-like geometric structures on $\mathbb{B}$ of a true quaternionic nature we set our sights on $\mathcal{M}(\mathbb{B})$, the regular Möbius transformations. Following the ideas from section 4 we first note that, in contrast to the arguments found in the proof of Theorem 5.1, the tensor $H_{0}$, and so $G_{0}$ and $\Omega_{0}$, is invariant under the regular Möbius transformations that fix the origin. We describe the latter, along with some properties, in the next result, whose first part follows from [12, Lemma 9.19].
@@
Let us denote by $\mathcal{M}(\mathbb{B})_{0}$ the set of regular Möbius transformations of $\mathbb{B}$ that fix 00. Then, the following properties hold.
@@
    yields maps of $\mathbb{B}$ that realize $\mathcal{M}(\mathbb{B})_{0}$. In other words, $\mathcal{M}(\mathbb{B})_{0}$ consists of the set of transformations $R_{u}$, for $u\in\mathrm{Sp}(1)$. In particular, $\mathcal{M}(\mathbb{B})_{0}$ is a group isomorphic to $\mathrm{Sp}(1)$ acting on the right on $\mathbb{B}$ and fixing 00.
@@
    The differential at 00 of the $\mathcal{M}(\mathbb{B})_{0}$-action on $\mathbb{B}$ leaves invariant the form $H_{0}$. In particular, the same holds for $G_{0}$ and $\Omega_{0}$.
@@
As noted above, (1) is a consequence of [12, Lemma 9.19].
Since the $\mathcal{M}_{0}(\mathbb{B})$-action on $\mathbb{B}$ is $\mathbb{R}$-linear, its differential has the same expression. Hence, the trivial identity
@@
The $\mathrm{Sp}(1)$-invariance of $H_{0}$ is the first step to build geometric structures on $\mathbb{B}$ out of regular Möbius transformations. It is important to note that Proposition 5.3 has provided a sort of isotropy at 00 for such transformations, which turns out to be an actual group. However, for the composition of maps, $\mathcal{M}(\mathbb{B})$ is not a group and so at this point the full group theoretic argument used in section 4 for the classical complex case breaks down. Nevertheless, the next result provides the analog of equation (4.4) for regular Möbius transformations. It is worthwhile to mention that the techniques from [5] are similar in spirit, since such work considers regular self-maps of $\mathbb{B}$ that fix the origin.
@@
[…]
@@
$H_{q}(\alpha,\beta)=\frac{(1-q^{2})^{-1}(\alpha-q\alpha q)(\overline{\beta-q\beta q})(1-\overline{q}^{2})^{-1}}{(1-|q|^{2})^{2}},$
@@
[…]
This proves that all three tensors $H$, $G$ and $\Omega$ are well defined. Hence, the definition of $H$ clearly implies claims (1) and (2).
For a given $q\in\mathbb{B}$, it is straightforward to compute the formula for $H_{q}$ by choosing $\mathcal{F}_{q}$ as the transformation mapping $q$ to 00 and using the expression for its differential at $q$ obtained in Proposition 3.3(2). The formula for $H$ now implies the smoothness of all three tensors.
Finally, to prove the positive definiteness of $H$ let us choose $q\in\mathbb{B}$ and $\alpha\in\mathbb{H}$. Then, we clearly have
@@ 6. Properties of the slice geometric structures on $\mathbb{B}$
The tensors $G$ and $\Omega$ can be given explicit expressions by simply taking the real and imaginary parts, respectively, of $H$. However, we can obtain simpler formulas. To achieve this, we introduce some notation.
@@
where $J$ is any element of $\mathbb{S}$ that anti-commutes with $I$. This decomposition is orthogonal for the canonical inner product of $\mathbb{H}\simeq\mathbb{R}^{4}$. In particular, $\mathbb{C}_{I}J$ is the orthogonal complement for such inner product and so it does not depend on the choice of $J$. We will denote from now on by $\pi_{I}$ and $\pi_{I}^{\perp}$ the orthogonal projections from $\mathbb{H}$ onto $\mathbb{C}_{I}$ and $(\mathbb{C}_{I})^{\perp}=\mathbb{C}_{I}J$, respectively.
It follows from the proof of Theorem 5.7 that $G$ is indeed a smooth Riemannian metric. The next result further provides for $G$ some useful simplified formulas. Recall that $\widehat{G}$ is the $\mathrm{Sp}(1,1)$-invariant metric on $\mathbb{B}$ from Theorem 5.1(1). Also, and as usual, $T_{q}\mathbb{B}$ will denote the tangent space to $\mathbb{B}$ at a point $q$; note that this space is naturally isomorphic (as a real vector space) to $\mathbb{H}$.
@@
$G_{q}(\alpha,\beta)$ $=\frac{\operatorname{Re}(\alpha\overline{\beta})}{(1-|q|^{2})^{2}}+4\frac{|\operatorname{Im}(q)|^{2}\operatorname{Re}(\pi_{I}^{\perp}(\alpha)\pi_{I}^{\perp}(\beta))}{|1-q^{2}|^{2}(1-|q|^{2})^{2}}$
$=\widehat{G}_{q}(\alpha,\beta)+4\frac{|\operatorname{Im}(q)|^{2}\operatorname{Re}(\pi_{I}^{\perp}(\alpha)\pi_{I}^{\perp}(\beta))}{|1-q^{2}|^{2}(1-|q|^{2})^{2}},$
for every $q\in\mathbb{D}_{I}$, where $I\in\mathbb{S}$, and $\alpha,\beta\in\mathbb{H}$. In particular, for every $\alpha\in T_{q}\mathbb{B}$, the length $\|\alpha\|_{q}$ of $\alpha$ with respect to the metric $G_{q}$ is given by
$\|\alpha\|_{q}^{2}=G_{q}(\alpha,\alpha)=\frac{|\alpha|^{2}}{(1-|q|^{2})^{2}}+4\frac{|\operatorname{Im}(q)|^{2}\operatorname{Re}(\pi_{I}^{\perp}(\alpha)^{2})}{|1-q^{2}|^{2}(1-|q|^{2})^{2}}$
for every $q\in\mathbb{C}_{I}$, where $I\in\mathbb{S}$.
@@
We will use in this proof the following easy to check fact
(6.1) $\operatorname{Re}(\alpha\beta)=\operatorname{Re}(\beta\alpha),$
for every $\alpha,\beta\in\mathbb{H}$. In particular, the first expression is now a consequence of this identity and Theorem 5.7 by the following computation
$\operatorname{Re}((1-q^{2})^{-1}$ $(\alpha-q\alpha q)(\overline{\beta-q\beta q})(1-\overline{q}^{2})^{-1})=$
$=\operatorname{Re}((1-\overline{q}^{2})^{-1}(1-q^{2})^{-1}(\alpha-q\alpha q)(\overline{\beta-q\beta q}))$
$=|1-q^{2}|^{-2}\operatorname{Re}((\alpha-q\alpha q)(\overline{\beta-q\beta q})),$
for every $q\in\mathbb{B}$ and $\alpha,\beta\in\mathbb{H}$.
We also note that for $q\in\mathbb{R}$, the second and third expressions for $G_{q}$ are obvious consequences of the first one, which was just proved. Hence, we proceed to manipulate the expression
@@
under the assumption that $q\in\mathbb{C}_{I}\setminus\mathbb{R}$ for some $I\in\mathbb{S}$. In particular, we can consider $\pi_{I}$ and $\pi_{I}^{\perp}$ and the decomposition that they entitle. We first compute as follows
$\operatorname{Re}((\alpha-q\alpha q)(\overline{\beta-q\beta q}))$ $=\operatorname{Re}(\alpha\overline{\beta})+\operatorname{Re}(q\alpha q\overline{q}\overline{\beta}\overline{q})-\operatorname{Re}(q\alpha q\overline{\beta})-\operatorname{Re}(\alpha\overline{q}\overline{\beta}\overline{q})$
$=\operatorname{Re}(\alpha\overline{\beta})+|q|^{2}\operatorname{Re}(q\alpha\overline{\beta}\overline{q})-\operatorname{Re}(q\alpha q\overline{\beta})-\operatorname{Re}(\overline{q}\alpha\overline{q}\overline{\beta})$
$=\operatorname{Re}(\alpha\overline{\beta})+|q|^{4}\operatorname{Re}(\alpha\overline{\beta})-\operatorname{Re}(q\alpha q\overline{\beta})-\operatorname{Re}(\overline{q}\alpha\overline{q}\overline{\beta})$
$=(1+|q|^{4})\operatorname{Re}(\alpha\overline{\beta})-\operatorname{Re}(q\alpha q\overline{\beta})-\operatorname{Re}(\overline{q}\alpha\overline{q}\overline{\beta}),$
where we have applied (6.1) on the second identity (fourth term) and third identity (second term). Let us choose $J\in\mathbb{S}$ that anti-commutes with $I$. If we write $\alpha=\alpha_{1}+\alpha_{2}J$ and $\beta=\beta_{1}+\beta_{2}J$, with $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}\in\mathbb{C}_{I}$, then a direct computation yields
@@
$=\big{(}1-2\operatorname{Re}(q^{2})+|q|^{4}\big{)}\operatorname{Re}(\alpha\overline{\beta})+2\big{(}\operatorname{Re}(q^{2})-|q|^{2}\big{)}\operatorname{Re}(\alpha_{2}\overline{\beta_{2}})$
$=|1-q^{2}|^{2}\operatorname{Re}(\alpha\overline{\beta})-4|\operatorname{Im}(q)|^{2}\operatorname{Re}(\alpha_{2}\overline{\beta_{2}}),$
@@
The third formula now follows from Theorem 5.1(1).
Finally, the formula for the length of a tangent vector with respect to the metric $G$ follows by taking $\alpha=\beta$ in the expressions already obtained. ∎
As a consequence, we conclude that the Riemannian geometry on the disk $\mathbb{D}_{I}=\mathbb{B}\cap\mathbb{C}_{I}$ induced from the slice Riemannian metric $G$ of $\mathbb{B}$ is exactly the usual hyperbolic geometry. More precisely, we have the next result. We will denote by $g_{I}$ the Riemannian metric on $\mathbb{D}_{I}$ induced by the canonical identification $\mathbb{D}_{I}\simeq\mathbb{D}$.